# A theoretical framework for self-supervised contrastive learning for continuous dependent data

### v2

@@ I Introduction
Self-supervised learning (SSL) methods demonstrate significant potential in achieving the quality of supervised models using a vast amount of unlabeled data [1, 2]. The SSL paradigm provides a method for training an encoder using a vast amount of unlabeled data, which produces a universal representation (embedding) of an input signal. Since representations are universal, we can use them to solve various downstream problem tasks without time- and data-consuming full retraining of the encoder [3].
@@ III-C3 Spatio-temporal objective function
III-D Further extensions
Similarly, one can account for other dependencies presented in the data. For example, seasonality imposes an additional component on the expected similarity between observations at different timestamps. Climate-related data is a clear, practical application with non-zero seasonal components, but similar patterns are also observed in retail and other datasets.
@@ IV-C1 Downstream problems description
Instead of solving the regression problem, i.e., predicting PDSI directly, we consider a binary classification task: whether it is a drought or not. To convert the regression task to a classification problem, we considered a case as a drought if the PDSI is less than 2-2 following [38].
@@ IV-C5 Results
V Conclusions

### v3

@@ I Introduction
Self-supervised learning (SSL) methods demonstrate significant potential in achieving the quality of supervised models using a vast amount of unlabeled data [1, 2, 3]. The SSL paradigm provides a method for training an encoder using a vast amount of unlabeled data, which produces a universal representation (embedding) of an input signal. Since representations are universal, we can use them to solve various downstream problem tasks without time- and data-consuming full retraining of the encoder [4].
@@ IV-C1 Downstream problems description
Instead of solving the regression problem, i.e., predicting PDSI directly, we consider a binary classification task: whether it is a drought or not. To convert the regression task to a classification problem, we considered a case as a drought if the PDSI is less than 2-2 following [39, 42].
@@ IV-C5 Results
V Discussion
The main assumption underlying our theoretical framework is data continuity. Nevertheless, the core idea can be naturally extended to other data domains. For graph-structured data, hard dependency corresponds to immediate node adjacency, while soft dependency can be defined through path lengths or spectral distances on the graph Laplacian. For sequential user data, recent actions form hard dependencies, whereas older interactions decay in relevance, leading to a soft autoregressive structure. For non-smooth or seasonal data, an additional seasonal component can be incorporated into the similarity matrix (analogous to SARIMA [45] decomposition), capturing periodic dependencies. Our approach can also handle anomalies by masking corrupted entries in the similarity matrix, thereby avoiding false positive pairs. Such a mask can be obtained either from external anomaly detection models (e.g., change-point detection) or directly inferred from domain-specific knowledge. Finally, in the presence of concept drift, the similarity matrix can be adaptively updated (e.g., in a sliding window fashion), which enables the framework to capture evolving dependency structures in streaming data. In summary, the theoretical foundation we propose is general and provides a principled way to extend self-supervised contrastive learning to a wide class of dependent data types.
VI Conclusions
@@
VII Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement with Skoltech №139-10-2025-033.